OpenStudy (anonymous):

Suppose 80 points are placed around a circle. A line segment is drawn between each pair of points.How many line segments are drawn?

OpenStudy (anonymous):

you have 80 points along a circle and you have to choose 2 of them to make a line segment the number of ways to do this is \(\binom{80}{2}\) some times written as \(_{80}C_2\) and read "80 choose 2"

OpenStudy (anonymous):

you know how to compute that number?

OpenStudy (anonymous):

I have not idea

OpenStudy (anonymous):

i take it this is some combinatorics class or maybe probability?

OpenStudy (anonymous):

it is my MGF1107 class LIberal Art II

OpenStudy (anonymous):

liberal arts?! lol have fun actually it is the same idea as your previous problem

OpenStudy (anonymous):

\[\binom{80}{2}=\frac{80\times 79}{2}\]

OpenStudy (anonymous):

yeah I am completely lost

OpenStudy (anonymous):

would the answer be 3160

OpenStudy (anonymous):

yeah i can see what you need to know is a formula for \(\binom{n}{k}\) the number of ways you can choose \(k\) items from a set of \(n\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the formula is \[\binom{n}{k}=\frac{n!}{k!(n-k)!}\] but you don't really use it exactly in this form have you seen this before?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh then i don't know how you are expected to do the problem without it what methods are you supposed to use?

OpenStudy (anonymous):

you certainly don't have time to count them all!

OpenStudy (anonymous):

I am not sure because this math class is online

OpenStudy (anonymous):

you got a book?

OpenStudy (anonymous):

yeah online and I am lost in the book

OpenStudy (anonymous):

I dont see the formula in the book at all

OpenStudy (anonymous):

hmm you need it for the last problem you asked as well, the maze problem

OpenStudy (anonymous):

maybe it is written as \(_nC_k\)

OpenStudy (anonymous):

is there anything there about pascal's triangle?

OpenStudy (anonymous):

I know i posted the maze i dont know if you can click on the link i posted

OpenStudy (anonymous):

let me see if i can find it, i can help you with that one as well

OpenStudy (anonymous):

thank you I sent you the link to my book

OpenStudy (anonymous):

yeah i got it what page are you on?

OpenStudy (anonymous):

problem solving

OpenStudy (anonymous):

do you know what page it is?

OpenStudy (anonymous):

page 13

OpenStudy (anonymous):

page 14

OpenStudy (anonymous):

jeez you are right, i guess you are just supposed to "figure it out"

OpenStudy (anonymous):

let me find your last problem and see if i can help you with it

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

I am so lost thank you for helping

OpenStudy (anonymous):

could you repost? i can't find the damned problem

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you can put it here if you like

OpenStudy (anonymous):

Morris Mouse can easily find his way through the maze from the entrance A to the exit B. However, he only receives food if he finds the exit without going west or south. (North is towards the top of the page.) How many different paths can he take through the maze to receive food? (Note: Different paths have at least one distinct section. See the diagram for an example.) https://angel.spcollege.edu/AngelUploads/QuestionData/2bfb73e4-236a-4599-bf10-af270b97f4c8/61324442322642544448.png# {7a7875da-913f-49bb-a105-e7c68ffc0aca}

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok now before we do this one lets redo the first one with a method that might be more suitable

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pick one of the 80 points it has 79 places to go pick the next one it has 78 place to go (because you don't want to count the first one twice) pick the next one it has 77 places to go etc

OpenStudy (anonymous):

so you can solve this by adding \[79+78+77+76+...+2+1\] or \[1+2+3+...+77+78+79\] now i do notice that you have a formula for this in the book

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the formula is \[1+2+3+...+n=\frac{n(n+1)}{2}\] which in your case is \[\frac{79\times 80}{2}\] same exact answer as before

OpenStudy (anonymous):

3160 was the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now for the next one, i cannot think of simple explanation but i can explain it

OpenStudy (anonymous):

i have 10 questions to do. Can you help me with them

OpenStudy (anonymous):

Thank you. Your such a big help

OpenStudy (anonymous):

lol sure until i get tired

OpenStudy (anonymous):

lol ok thank you. I need to pass this class for my RN degree

OpenStudy (anonymous):

in the maze problem you have, you can think of it this way: you have to go to the right 5 times and up 5 times, in some order

OpenStudy (anonymous):

ok I am looking at it now

OpenStudy (anonymous):

let me know when it is clear that you are going to take 5 steps right and 5 steps up to get from point A to point B

OpenStudy (anonymous):

ok so I took 5 steps to the right and 5 up

OpenStudy (anonymous):

that would give me 25 steps

OpenStudy (anonymous):

now the question is, how many ways can you do that here is one way \[(r, r, u, r, u, u, r, u, r, u)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if i replace all \(r\) by \(0\) and all \(u\) by \(1\) we turn \[(r, r, u, r, u, u, r, u, r, u)\] in to \[(0,0,1,0,1,1,0,1,0,1)\]

OpenStudy (anonymous):

now I am lost

OpenStudy (anonymous):

forget that then, it is unnecessary

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

the question is, how many ways can i fill ten slots with 5 \(r\) and 5 \(u\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

now we have to use the formula i wrote above i see no way around it

OpenStudy (anonymous):

ok let me right that formula down

OpenStudy (anonymous):

the question translates in to this "how many ways can i pick 5 out of the 10 slots to put an \(r\)?"

OpenStudy (anonymous):

and the answer is \[\binom{10}{5}=\frac{10\times 9\times 8\times 7\times 6}{5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

as i said this i called "10 choose 5" and we can ask wolfram for the answer, although this is very easy to compute if you cancel first and multiply last

OpenStudy (anonymous):

ok let me solve this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

252 is the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yw next?

OpenStudy (anonymous):

What is the 76th decimal digit in the decimal representation of

OpenStudy (anonymous):

13/11

OpenStudy (anonymous):

that can't be that hard, you only have two to choose from

OpenStudy (anonymous):

What is the 76th decimal digit in the decimal representation of 13/11 https://angel.spcollege.edu/AngelUploads/QuestionData/acae6292-47fd-4c42-9c1a-f19417bc9091/62K31R313554314L4653.png# {2752460f-f064-4665-9c63-2331704bfd48}

OpenStudy (anonymous):

76 is an even number right?

OpenStudy (anonymous):

so it is 8

OpenStudy (anonymous):

lots easier than the last two by far

OpenStudy (anonymous):

yes that is what i got

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok i hope the rest are that easy

OpenStudy (anonymous):

If two ladders are placed end to end, their combined height is 46.5 feet.One ladder is 3.5 feet shorter than the other ladder.What are the heights of the two ladders?

OpenStudy (anonymous):

we are on number 5

OpenStudy (anonymous):

you got lots of way to do this you can write an equation if you like or just guess and check

OpenStudy (anonymous):

ok what way would be the easy

OpenStudy (anonymous):

i am always accused of doing this the not math way but if it was me i would guess if one is 20 and the other is 3.5 longer then it would be 23.5 but then the total would only be \(20+23.5=43.5\) so that is not the right answer try a bigger one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if you want an equation, call one length \(x\) so the other would be \(x+3.5\) and then solve \[x+x+3.5=46.5\]

OpenStudy (anonymous):

or as we say in math, \[2x+3.5=46.5\]

OpenStudy (anonymous):

can you solve that one? takes two steps subtract \(3.5\) from both sides, divide both sides by \(2\)

OpenStudy (anonymous):

ok 21.5

OpenStudy (anonymous):

answer would be 21.5 feet and 18 feet